# TPP-LLM: Modeling Temporal Point Processes by Efficiently Fine-Tuning Large Language Models

### v1

@@ 1 Introduction
With the rise of powerful large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and Llama-3 (Dubey et al., 2024), new opportunities have emerged for using LLMs to understand and predict event sequences by capturing rich semantic and contextual information. Inspired by their success in text-based tasks (Zhao et al., 2023) and time series prediction (Zhou et al., 2023; Jin et al., 2023a; Zhang et al., 2024b), we propose TPP-LLM [^1] GitHub repository: https://github.com/zefang-liu/TPP-LLM (Figure 1), a novel framework that integrates LLMs with TPPs to model both the temporal and semantic aspects of event sequences. By leveraging pretrained LLMs, TPP-LLM directly utilizes textual descriptions of event types, moving beyond traditional methods that rely on categorical representations. To ensure the model captures temporal dynamics, we incorporate temporal embeddings alongside these event descriptions. To efficiently adapt LLMs for TPP modeling, we employ low-rank adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning (PEFT) (Liu et al., 2022) method, allowing us to adjust a small subset of LLM parameters, reducing computational cost while maintaining high performance. Through extensive experiments on real-world datasets, we demonstrate that TPP-LLM consistently outperforms state-of-the-art baselines in sequence modeling and next event prediction.
@@ 2 Related Work
Neural Temporal Point Processes. Recent advances in neural temporal point processes (TPPs) have introduced models that leverage deep learning techniques to capture complex temporal dependencies and event interactions. Many of these models use recurrent neural networks (RNNs) (Hochreiter, 1997) or self-attention mechanisms (Vaswani et al., 2017) to model event intensities based on event history. For example, RMTPP (Du et al., 2016) and NHP (Mei & Eisner, 2017) use RNNs to learn temporal influences, while more recent approaches like SAHP (Zhang et al., 2020) and THP (Zuo et al., 2020) utilize self-attention to capture long-term dependencies. Other models, such as those based on neural ODEs (Chen et al., 2020), normalizing flows (Shchur et al., 2019), attention mechanisms (Yang et al., 2022), diffusion processes (Yuan et al., 2023), meta learning (Bae et al., 2023), and Mamba models (Gao et al., 2024), offer flexible and high-fidelity modeling of discrete events in continuous time. These methods have significantly improved the performance of TPPs by modeling complex interactions and dynamic event relationships.
Large Language Models for Event Sequences. Recent work has explored integrating large language models (LLMs) into event sequence prediction tasks (Jin et al., 2023b). Shi et al. (2024) propose LAMP, a framework that leverages LLMs for abductive reasoning to improve event sequence prediction. Xue et al. (2023b) introduce PromptTPP, which incorporates continual learning into neural temporal point processes to enable adaptive and efficient learning of streaming event sequences. Song et al. (2024) present LaTee, a model utilizing an amortized expectation-maximization framework with logic trees as latent variables and a learnable GFlowNet to generate logic tree samples for more effective event reasoning.
@@ 3.1 Temporal Point Processes
$\lambda(t,k|\mathcal{H}_{t})=\lim_{\Delta t\to 0}\frac{\mathbb{E}[N(t+\Delta t)-N(t)|\mathcal{H}_{t}]}{\Delta t},$ (1)
where $\mathcal{H}_{t}=\{(t_{j},k_{j}):t_{j}<t\}$ represents the history of previous events up to time $t$, and $N(t)$ is the counting process representing the number of events that have occurred up to time $t$. This intensity function provides the expected number of events occurring in a small time interval $[t,t+\Delta t)$, conditioned on the past. The joint probability density $p(t,k|\mathcal{H}_{t})$ represents the likelihood of the next event occurring at time $t$ with type $k$, conditioned on the history $\mathcal{H}_{t}$. It is expressed as: $p(t,k|\mathcal{H}_{t})=\lambda(t,k|\mathcal{H}_{t})\exp\left(-\int_{t_{i}}^{t}\sum_{k^{\prime}\in\mathcal{K}}\lambda(s,k^{\prime}|\mathcal{H}_{s})\,\mathrm{d}s\right)$, where the integral accounts for no events occurring between the last event at $t_{i}$ and the current time $t$, capturing both event timing and type dependencies.
@@ 3.2 Neural Temporal Point Processes
In neural TPPs, for each event $(t_{i},k_{i})$, an embedding $\bm{e}_{i}\in\mathbb{R}^{D}$ is computed through embedding layers based on the event time $t_{i}$ and the event type $k_{i}$. The hidden state $\bm{h}_{t}$, which represents the history up to time $t$, is then updated based on the current event’s embedding and the previous hidden state $\bm{h}_{i-1}$. This update can be formulated as: $\bm{h}_{i}=f_{\text{update}}(\bm{h}_{i-1},\bm{e}_{i})$, where $f_{\text{update}}$ is a neural network, often implemented as a recurrent neural network (RNN) (Hochreiter, 1997) or a more advanced attention-based mechanism (Vaswani et al., 2017). With the updated hidden state $\bm{h}_{i}$, the next event time $t_{i+1}$ and event type $k_{i+1}$ are sampled from the probability distribution conditioned on $\bm{h}_{i}$: $t_{i+1},k_{i+1}\sim P(t_{i+1},k_{i+1}|\bm{h}_{i})$. Different neural TPP models employ various architectures for the state update function $f$. Early approaches (Du et al., 2016; Mei & Eisner, 2017) use RNNs to capture temporal dependencies between events, while more recent models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) replace the recurrent structure with attention-based layers, allowing for better long-range interactions. These neural-based methods enhance the flexibility of TPPs, learning event dependencies from complex datasets in a data-driven manner.
@@ 4.1 Event and Prompt Embeddings
$[\bm{t}_{i}]_{j}=\begin{cases}\cos\left(\frac{t_{j}}{10000^{(i-1)/D}}\right),&\text{when }i\text{ is odd},\\ \sin\left(\frac{t_{j}}{10000^{i/D}}\right),&\text{when }i\text{ is even}.\end{cases}$ (3)
Other encoding methods (Zhang et al., 2020; Gao & Dai, 2024) can also be applied.
@@ 4.2 History Vectors and Intensity Function
The entire sequence $\bm{X}$ is then passed through the LLM to obtain contextualized hidden states for each token: $\bm{H}=\text{LLM}(\bm{X})$. After processing, we extract the hidden states corresponding to the last embedding vector of each event. For example, the hidden state of the event $i$ is $\bm{h}_{i}=\bm{H}_{L_{P}+\sum_{j\leq i}L_{j}+i}\in\mathbb{R}^{H}$. The selected hidden states $\bm{h}_{i}$ represent the event history up to time $t_{i}$ (inclusive): $\mathcal{H}_{t_{i}}^{\prime}=\{(t_{j},k_{j}):t_{j}\leq t_{i}\}$. These history vectors are then used for modeling TPPs.
In our model, the intensity function is parameterized using the history vector $\bm{h}_{i}$, which encodes the event history from the initial time to time $t_{i}$. To compute the intensity function between $t_{i}$ and $t_{i+1}$, we apply the linear transformation to the hidden state $\bm{h}_{i}$. For the event type $k$, the intensity function (Zuo et al., 2020; Gao & Dai, 2024) is modeled as:
@@ 5.1 Datasets
We conduct experiments on five real-world datasets: Stack Overflow, Chicago Crime, NYC Taxi Trip, U.S. Earthquake, and Amazon Review. Their statistics are shown in Table 1. The datasets span various applications and are widely used in prior TPP research, making them well-suited for evaluating the performance of our model. However, since the currently available versions lack the corresponding event type texts required by TPP-LLM, we preprocess data to include these critical textual descriptions. These diverse datasets allow us to assess the model’s generalization capabilities across different domains, handling sequences with varying lengths, event types, and temporal resolutions. More detailed information is available in Appendix A.
Table 1: Datasets statistics overview for event sequences.
@@ 5.2 Prompt Design
We design the prompt to provide a structured guide for the model, helping it understand the task and the sequence of events effectively. The prompt includes essential details such as the sequence context and specifics about event types, allowing the model to focus on the key components it needs to process for accurate predictions. The general structure of the prompt is as follows: “{Sequence Description} {Event Description} {Task Description}” with the task description tailored to the prediction task. When event type precedes time in the embedding sequences, the task is framed as: “Based on this sequence, predict the next event type and the corresponding time.” Alternatively, when event time comes first, the task becomes: “Based on this sequence, predict the next event time and the corresponding type.” The specific sequence and event descriptions for datasets used in our experiments are listed in Appendix C.
@@ 5.3 Baselines and Evaluation Metrics
We compare our model, TPP-LLM, with several state-of-the-art (SOTA) baselines to evaluate its performance across different tasks. The baselines include the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), Self-Attentive Hawkes Process (SAHP) (Zhang et al., 2020), Transformer Hawkes Process (THP) (Zuo et al., 2020), and Attentive Neural Hawkes Process (AttNHP) (Yang et al., 2022). These models represent leading approaches in neural TPP modeling. Detailed descriptions of the baselines are provided in Appendix B.
@@ 5.4 Experimental Setup
We conduct experiments using two foundation models for TPP-LLM: TinyLlama-1.1B-Chat-v1.0 (Zhang et al., 2024a) and Gemma-2-2B-IT (Team et al., 2024), both of which are quantized to 4-bit precision (Dettmers et al., 2024) for efficient GPU memory usage. To capture temporal dynamics, we use temporal positional encoding (Zuo et al., 2020), and event type embeddings are processed first, followed by the temporal embedding for each event. The non-event integral term in the log-likelihood is handled using Monte Carlo integration (Zuo et al., 2020) with 20 samples per time interval, applied consistently across all models. For fine-tuning, we employ LoRA (Hu et al., 2021) by adapting weight matrices in attention modules, with dropout applied but without bias. The Adam optimizer (Kingma, 2014) is used for optimizing both the LoRA layers and prediction layers. Baselines implemented in the EasyTPP framework (Xue et al., 2023a) are utilized, with hyperparameters adapted from it for a fair comparison. Experiments results are averaged over five runs with early stopping, and additional hyperparameters are provided in Appendix D. We utilize a single NVIDIA A10 or A100 GPU for baselines and a single H100 GPU for TPP-LLM.
5.5 Experiment Results
@@
Figure 2: Performance comparison of next event type prediction accuracy across five different datasets. Each subplot shows the accuracy of six models with error bars.
@@
Figure 3: Performance comparison of next event time prediction RMSE across five different datasets. Each subplot shows the RMSE of six models with error bars.
@@
5.6 Few-Shot Learning
Model StackOverflow Crime Taxi Earthquake Amazon
NHP -6.515 -6.202 -1.236 -0.697 -4.106
SAHP -6.839 -6.301 -1.173 -0.615 -4.567
THP -2.683 -2.926 -0.878 -0.625 -1.768
AttNHP -2.070 -2.594 -0.670 -0.521 -1.474
TPP-Llama -2.026 -2.668 -0.098 -0.539 -1.469
TPP-Gemma -2.068 -2.716 0.051 -0.600 -1.630
Model StackOverflow Crime Taxi Earthquake Amazon
@@ 5.7 Ablation Studies
To understand the contribution of different components in TPP-LLM, we conduct a series of ablation studies. By systematically removing or altering key parts of the model, we analyze how each element affects overall performance and identify which configurations lead to the best results.
@@ 5.7.1 Foundation Models
The performance comparison in Table 6 shows the impact of different LLMs on TPP-LLM’s performance. TinyLlama-1.1B-Chat-v1.0 and TinyLlama-1.1B-Intermediate show similar log-likelihood and accuracy scores, with Chat slightly outperforming on next event type prediction for Stack Overflow and U.S. Earthquake. Gemma-2-2B-IT achieves the best RMSE for event time prediction on NYC Taxi Trip and U.S. Earthquake, highlighting its strength in modeling temporal dynamics. Overall, both TinyLlama and Gemma perform competitively, with each excelling in different metrics. The consistent performance across models underscores the robustness of TPP-LLM.
@@ 5.7.2 Temporal Embeddings
As shown in Table 7, the choice of embedding type and order affects model performance. Positional encoding generally outperforms linear embedding across most metrics. When event time embeddings are processed first, positional encoding achieves the best next event type prediction accuracy and competitive RMSE values. Linear embedding also performs well, with the best log-likelihood on U.S. Earthquake when event time comes first. The results suggest that processing event time before type improves event type prediction, while placing time last enhances time prediction accuracy. These findings highlight the importance of embedding order in model design.
@@
Embedding First StackOverflow Earthquake
Positional Event Type -1.$7$0.475/62.70%/0.288
Positional Event Time -1.$7$0.475/63.15%/0.307
Linear Event Type -1.768/44.15%/0.478 -0.474/62.65%/0.297
Linear Event Time -1.768/44.20%/0.477 -0.467/63.01%/0.343
5.7.3 Prompt Settings
Foundation Model Prompt StackOverflow Earthquake
TinyLlama-1.1B-Intermediate Y -1.$7$0.476/62.65%/0.299
TinyLlama-1.1B-Chat-v1.0 Y -1.$7$0.475/62.70%/0.288
5.7.4 Fine-Tuning Settings
Table 9 and Figure 4 illustrate the impact of different LoRA rank configurations on performance. Without fine-tuning (only training the head layers), the model suffers significant drops in log-likelihood and accuracy, highlighting the need for adapting the pretrained LLM. Higher LoRA ranks generally improve results, with rank 32 achieving the best accuracy and log-likelihood on U.S. Earthquake, while lower ranks perform well in terms of log-likelihood and event time prediction. Fine-tuning with LoRA consistently enhances performance, with higher ranks benefiting more complex tasks and lower ranks offering competitive results with less computational cost.
Table 9: Performance comparison of log-likelihood, next event type prediction accuracy, and next event time prediction RMSE across different datasets with various fine-tuning settings.
Fine-Tuning StackOverflow Earthquake
None -1.891/42.43%/0.484 -0.497/62.95%/0.306
LoRA (rank 4) -1.774/44.18%/0.474 -0.486/62.78%/0.291
LoRA (rank 8) -1.767/44.20%/0.484 -0.480/63.19%/0.296
LoRA (rank 16) -1.$7$0.475/62.70%/0.288
LoRA (rank 32) -1.771/44.39%/0.482 -0.475/63.24%/0.292
(a) Log-likelihood
(b) Accuracy
(c) RMSE

### v2

@@ 1 Introduction
With the rise of powerful large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and Llama-3 (Dubey et al., 2024), new opportunities have emerged for using LLMs to understand and predict event sequences by capturing rich semantic and contextual information. Inspired by their success in text-based tasks (Zhao et al., 2023) and time series prediction (Zhou et al., 2023; Jin et al., 2023a; Zhang et al., 2024b), we propose TPP-LLM [^1] GitHub repository available on https://github.com/zefang-liu/TPP-LLM. (Figure 1), a novel framework that integrates LLMs with TPPs to model both the temporal and semantic aspects of event sequences. By leveraging pretrained LLMs, TPP-LLM directly utilizes textual descriptions of event types, moving beyond traditional methods that rely on categorical representations. To ensure the model captures temporal dynamics, we incorporate temporal embeddings alongside these event descriptions. To efficiently adapt LLMs for TPP modeling, we employ low-rank adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning (PEFT) (Liu et al., 2022) method, allowing us to adjust a small subset of LLM parameters, reducing computational cost while maintaining high performance. Through extensive experiments on real-world datasets, we demonstrate that TPP-LLM consistently outperforms state-of-the-art baselines in sequence modeling and next event prediction.
@@ 2 Related Work
Neural Temporal Point Processes. Recent advances in neural temporal point processes (TPPs) have introduced models that leverage deep learning techniques to capture complex temporal dependencies and event interactions. Many of these models use recurrent neural networks (RNNs) (Hochreiter et al., 1997) or self-attention mechanisms (Vaswani et al., 2017) to model event intensities based on event history. For example, RMTPP (Du et al., 2016) and NHP (Mei & Eisner, 2017) use RNNs to learn temporal influences, while more recent approaches like SAHP (Zhang et al., 2020) and THP (Zuo et al., 2020) utilize self-attention to capture long-term dependencies. Other models, such as those based on fully neural networks (Omi et al., 2019), normalizing flows (Shchur et al., 2019), neural ordinary differential equations (ODEs) (Chen et al., 2020), attention mechanisms (Yang et al., 2022), diffusion processes (Yuan et al., 2023), meta learning (Bae et al., 2023), and Mamba models (Gao et al., 2024), offer flexible and high-fidelity modeling of discrete events in continuous time. These methods have significantly improved the performance of TPPs by modeling complex interactions and dynamic event relationships.
Large Language Models for Event Sequences. Recent work has explored integrating large language models (LLMs) into event sequence prediction tasks (Jin et al., 2023b). Shi et al. (2024) propose LAMP, a framework that leverages LLMs for abductive reasoning to improve event sequence prediction. Xue et al. (2023) introduce PromptTPP, which incorporates continual learning into neural temporal point processes to enable adaptive and efficient learning of streaming event sequences. Song et al. (2024) present LaTee, a model utilizing an amortized expectation-maximization framework with logic trees as latent variables and a learnable GFlowNet to generate logic tree samples for more effective event reasoning.
@@ 3.1 Temporal Point Processes
$\lambda(t,k|\mathcal{H}_{t})=\lim_{\Delta t\to 0}\frac{\mathbb{E}[N_{k}(t+\Delta t)-N_{k}(t)|\mathcal{H}_{t}]}{\Delta t},$ (1)
where $\mathcal{H}_{t}=\{(t_{j},k_{j}):t_{j}<t\}$ represents the history of previous events up to time $t$, and $N_{k}(t)$ is the counting process representing the number of events that have occurred up to time $t$. This intensity function provides the expected number of events occurring in a small time interval $[t,t+\Delta t)$, conditioned on the past. The joint probability density $p(t,k|\mathcal{H}_{t})$ represents the likelihood of the next event occurring at time $t$ with type $k$, conditioned on the history $\mathcal{H}_{t}$. It is expressed as: $p(t,k|\mathcal{H}_{t})=\lambda(t,k|\mathcal{H}_{t})\exp\left(-\int_{t_{i}}^{t}\sum_{k^{\prime}\in\mathcal{K}}\lambda(s,k^{\prime}|\mathcal{H}_{s})\,\mathrm{d}s\right)$, where the integral accounts for no events occurring between the last event at $t_{i}$ and the current time $t$, capturing both event timing and type dependencies.
@@ 3.2 Neural Temporal Point Processes
In neural TPPs, for each event $(t_{i},k_{i})$, an embedding $\bm{e}_{i}\in\mathbb{R}^{D}$ is computed through embedding layers based on the event time $t_{i}$ and the event type $k_{i}$. The hidden state $\bm{h}_{t}$, which represents the history up to time $t$, is then updated based on the current event’s embedding and the previous hidden state $\bm{h}_{i-1}$. This update can be formulated as: $\bm{h}_{i}=f_{\text{update}}(\bm{h}_{i-1},\bm{e}_{i})$, where $f_{\text{update}}$ is a neural network, often implemented as a recurrent neural network (RNN) (Hochreiter et al., 1997) or a more advanced attention-based mechanism (Vaswani et al., 2017). With the updated hidden state $\bm{h}_{i}$, the next event time $t_{i+1}$ and event type $k_{i+1}$ are sampled from the probability distribution conditioned on $\bm{h}_{i}$: $t_{i+1},k_{i+1}\sim P(t_{i+1},k_{i+1}|\bm{h}_{i})$. Different neural TPP models employ various architectures for the state update function $f$. Early approaches (Du et al., 2016; Mei & Eisner, 2017) use RNNs to capture temporal dependencies between events, while more recent models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) replace the recurrent structure with attention-based layers, allowing for better long-range interactions. These neural-based methods enhance the flexibility of TPPs, learning event dependencies from complex datasets in a data-driven manner.
@@ 4.1 Event and Prompt Embeddings
$[\bm{t}_{i}]_{j}=\begin{cases}\cos\left(\frac{t_{i}}{10000^{(j-1)/D}}\right),&\text{when }j\text{ is odd},\\ \sin\left(\frac{t_{i}}{10000^{j/D}}\right),&\text{when }j\text{ is even}.\end{cases}$ (3)
Other temporal encoding methods (Zhang et al., 2020; Gao & Dai, 2024) can also be applied.
@@ 4.2 History Vectors and Intensity Function
The entire sequence $\bm{X}$ is then passed through the decoder-only transformer (LLM) to obtain contextualized hidden states for each token: $\bm{H}=\text{LLM}(\bm{X})$. After processing, we extract the hidden states corresponding to the last embedding vector of each event. For example, the hidden state of event $i$ is $\bm{h}_{i}=\bm{H}_{L_{P}+\sum_{j\leq i}L_{j}+i}\in\mathbb{R}^{H}$. The selected hidden state $\bm{h}_{i}$ represents the event history up to time $t_{i}$: $\mathcal{H}_{t_{i}}^{\prime}=\{(t_{j},k_{j}):t_{j}\leq t_{i}\}$. These history vectors are then used for modeling TPPs.
In our model, the intensity function is parameterized using the history vector $\bm{h}_{i}$, which encodes the event history from the initial time to time $t_{i}$. To compute the intensity function between $t_{i}$ and $t_{i+1}$, we apply the linear transformation to the hidden state $\bm{h}_{i}$. For the event type $k$, the intensity function (Du et al., 2016; Zuo et al., 2020; Gao & Dai, 2024) is modeled as:
@@ 5.1 Datasets
We conduct experiments on five real-world datasets [^2] Datasets available on https://huggingface.co/tppllm.: Stack Overflow, Chicago Crime, NYC Taxi Trip, U.S. Earthquake, and Amazon Review. Their statistics are shown in Table 1. The datasets span various applications and are widely used in prior TPP research, making them well-suited for evaluating the performance of our model. However, since the currently available versions lack the corresponding event type texts required by TPP-LLM, we preprocess data to include these critical textual descriptions. These diverse datasets allow us to assess the model’s generalization capabilities across different domains, handling sequences with varying lengths, event types, and temporal resolutions. More detailed information is available in Appendix A.
Table 1: Dataset statistics overview for event sequences.
@@ 5.2 Prompt Design
We design the prompt to provide a structured guide for the model, helping it understand the task and the sequence of events effectively. The prompt includes essential details such as the sequence context and specifics about event types, allowing the model to focus on the key components it needs to process for accurate predictions. The general structure of the prompt is as follows: “{Sequence Description} {Event Description} {Task Description}” with the task description tailored to the prediction task. When event type precedes time in the embedding sequences, the task is framed as: “Based on this sequence, predict the next event type and the corresponding time.” Alternatively, when event time comes first, the task becomes: “Based on this sequence, predict the next event time and the corresponding type.” The specific sequence and event descriptions for datasets used in our experiments are listed in Appendix B.
@@ 5.3 Baselines and Evaluation Metrics
We compare our model, TPP-LLM, with several state-of-the-art (SOTA) baselines to evaluate its performance across different tasks. The baselines include the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), Self-Attentive Hawkes Process (SAHP) (Zhang et al., 2020), Transformer Hawkes Process (THP) (Zuo et al., 2020), Attentive Neural Hawkes Process (AttNHP) (Yang et al., 2022), and Neural ODE-based Temporal Point Process (ODETPP) (Chen et al., 2020). These models represent leading approaches in neural TPP modeling. Detailed descriptions of the baselines are provided in Appendix C.
@@ 5.4 Experimental Setup
We conduct experiments using two foundation models for TPP-LLM: TinyLlama-1.1B-Chat-v1.0 (Zhang et al., 2024a) and Gemma-2-2B-IT (Team et al., 2024), both of which are quantized to 4-bit precision (Dettmers et al., 2024) for efficient GPU memory usage. To capture temporal dynamics, we use temporal positional encoding (Zuo et al., 2020), and event type embeddings are processed first, followed by the temporal embedding for each event. The non-event integral term in the log-likelihood is handled using Monte Carlo integration (Zuo et al., 2020) with 20 samples per time interval, applied consistently across all models. For fine-tuning, we employ LoRA (Hu et al., 2021) by adapting weight matrices in attention modules, with dropout applied but without bias. The Adam optimizer (Kingma, 2014) is used for optimizing both the LoRA layers and prediction layers. Baselines implemented in the EasyTPP framework (Xue et al., 2024) are utilized, with hyperparameters adapted from it for a fair comparison. Experiments results are averaged over five runs with early stopping, and additional hyperparameters and setup are provided in Appendix D and E. We utilize a single NVIDIA A10 and A100 GPU for baselines and a single H100 GPU for TPP-LLM.
5.5 Experimental Results
@@
Figure 2: Performance comparison of next event type prediction accuracy across five different datasets. Each subplot shows the accuracy of models with error bars.
@@
Figure 3: Performance comparison of next event time prediction RMSE across five different datasets. Each subplot shows the RMSE of models with error bars.
@@ 5.6 Ablation Studies
To understand the contribution of different components in TPP-LLM, we conduct a series of ablation studies. By systematically removing or altering key parts of the model, we analyze how each element affects overall performance and identify which configurations lead to the best results. More ablation stuides can be found in Appendix G.
@@ 5.6.1 Foundation Models
The performance comparison in Table 4 shows the impact of different LLMs on TPP-LLM’s performance. TinyLlama-1.1B-Chat-v1.0 and TinyLlama-1.1B-Intermediate show similar log-likelihood and accuracy scores, with Chat slightly outperforming on next event type prediction for Stack Overflow and U.S. Earthquake. Gemma-2-2B-IT achieves the best RMSE for event time prediction on NYC Taxi Trip and U.S. Earthquake, highlighting its strength in modeling temporal dynamics. The Llama-3.2 models (Dubey et al., 2024) excel in log-likelihood for Stack Overflow and U.S. Earthquake, with Llama-3.2-1B-Instruct achieving the highest accuracy for NYC Taxi Trip and U.S. Earthquake, showcasing their strong performance across diverse metrics and datasets. Overall, the consistent performance across models underscores the robustness of TPP-LLM.
@@
Llama-3.2-1B -1.770/44.12%/0.486 0.318/91.75%/0.875 -0.472/62.86%/0.301
Llama-3.2-1B-Instruct -1.775/44.15%/0.483 0.310/91.78%/0.886 -0.481/63.14%/0.303
Llama-3.2-3B -1.777/44.16%/0.491 0.285/91.69%/0.921 -0.466/62.84%/0.312
Llama-3.2-3B-Instruct -1.771/44.19%/0.494 0.300/91.57%/0.883 -0.469/63.02%/0.311
@@ 5.6.2 Temporal Embeddings
As shown in Table 5, the type and order of temporal embeddings influence model performance. Temporal positional encoding generally outperforms both time-shifted positional encoding and linear embeddings in most cases. Specifically, when event time embeddings are processed first, temporal positional encoding yields the best next event type prediction accuracy and competitive RMSE values on Stack Overflow and U.S. Earthquake. Linear embeddings also show strong results, with the best log-likelihood on U.S. Earthquake when event time is placed first. Time-shifted positional encoding exhibits lower performance across all metrics. These findings suggest that processing the event time before the event type improves event type prediction, while adjusting the embedding strategy can optimize model performance for different metrics.
@@
Temporal Embedding First Embedding StackOverflow Earthquake
Temporal Positional Encoding Event Type -1.$7$0.475/62.70%/0.288
Temporal Positional Encoding Event Time -1.$7$0.475/63.15%/0.307
Time-Shifted Positional Encoding Event Type -1.782/44.13%/0.550 -0.475/62.97%/0.360
Time-Shifted Positional Encoding Event Time -1.812/43.99%/0.556 -0.474/63.11%/0.385
Linear Layer Event Type -1.768/44.15%/0.478 -0.474/62.65%/0.297
Linear Layer Event Time -1.768/44.20%/0.477 -0.467/63.01%/0.343
5.7 Intensity Functions
We conduct an ablation study to compare three intensity functions adapted for TPP-Llama, all leveraging $\bm{h}i$ as the history vector. The modified THP intensity function (Zuo et al., 2020), $\text{softplus}(\alpha_{k}(t-t_{i})+\bm{w}{k}^{\mathsf{T}}\bm{h}i+b_{k})$, achieves the best overall performance, balancing flexibility and stability in capturing temporal dynamics. The RMTPP intensity function (Du et al., 2016), $\exp(\alpha_{k}(t-t_{i})+\bm{w}{k}^{\mathsf{T}}\bm{h}i+b_{k})$, performs well in event type prediction but slightly underperforms in event time prediction due to its exponential nature. The SAHP intensity function (Zhang et al., 2020), $\text{softplus}(\mu_{i}+(\eta_{i}-\mu_{i})\exp(-\gamma_{i}(t-t_{i})))$, where $\mu{i}=\text{gelu}(\bm{W}{\mu}\bm{h}{i})$, $\eta_{i}=\text{gelu}(\bm{W}{\eta}\bm{h}{i})$, and $\gamma_{i}=\text{gelu}(\bm{W}{\gamma}\bm{h}{i})$, shows lower performance across most metrics. These results highlight TPP-Llama’s robustness with different intensity functions while indicating that the modified THP intensity function provides the most consistent and reliable results for capturing temporal patterns effectively.
Table 6: Performance comparison of log-likelihood, accuracy, and RMSE using different intensity functions on various datasets.
Model Intensity Function StackOverflow Earthquake
TPP-Llama THP -1.$7$0.475/62.70%/0.288
TPP-Llama RMTPP -1.786/44.32%/0.484 -0.488/63.10%/0.361
TPP-Llama SAHP -4.689/42.21%/0.486 -0.548/62.88%/0.308